OpenStudy (anonymous):

Which of the following functions has the function rule y = x + 4? A.{(0, 2), (-2, -6), (1, 5)} B.{(2, 6), (-3, -7), (0, 4)} C.{(-3, 1), (0, 4), (2, 6)} D.{(-2, 2), (-1, -5), (3, 7)}

OpenStudy (astrophysics):

Try the solutions they gave you and see if left side = right side.

OpenStudy (astrophysics):

Look at b, you can see y = x+ 4, so plug 2 in for x, and 6 for y, 6=2+4 6 = 6 so that's right, now do it with the other ones.

OpenStudy (anonymous):

sooo.... C?

OpenStudy (astrophysics):

Yeah :)

OpenStudy (anonymous):

thanks